OpenStudy (epicteatime):

If 143 grams of chromium react with an excess of oxygen, as shown in the balanced chemical equation below, how many grams of chromium oxide can be formed? 4Cr + 3O2 -> 2Cr2O3

OpenStudy (epicteatime):

@phi

OpenStudy (epicteatime):

@shamil98

OpenStudy (epicteatime):

@ParthKohli

OpenStudy (epicteatime):

@Preetha

OpenStudy (epicteatime):

@petiteme

OpenStudy (kc_kennylau):

(a) Find the number of moles of chromium. (b) Find the number of moles of chromium oxide. (c) Find the number of grams of chromium oxide.

OpenStudy (epicteatime):

ok thanks

OpenStudy (kc_kennylau):

no problem

OpenStudy (epicteatime):

but how do i find the gram of CrO?

OpenStudy (kc_kennylau):

No you don't, it's asking for the gram of Cr2O3

OpenStudy (epicteatime):

ok

OpenStudy (epicteatime):

i got the 2 numbers, what do i do next? 103.992 47.997

OpenStudy (kc_kennylau):

What numbers are these?

OpenStudy (epicteatime):

the numbers after i multiplied them. So 103.992 is Cr's mole number x 2 And 47.997 is 15.999 x 3

OpenStudy (kc_kennylau):

Multiplying Cr's mole number x 2 doesn't get you anything -_-

OpenStudy (epicteatime):

;(

OpenStudy (epicteatime):

then what should i do afterwards?

OpenStudy (kc_kennylau):

convert 143 grams of chromium to moles

OpenStudy (epicteatime):

oh!!

OpenStudy (kc_kennylau):

yep :)

OpenStudy (epicteatime):

143 x 1/255.981= 0.559?

OpenStudy (epicteatime):

i dont know

OpenStudy (epicteatime):

@JFraser

OpenStudy (jfraser):

what is that 0.559 that you have? what UNIT should it have?

OpenStudy (epicteatime):

so is it correct, other than the units missing?

OpenStudy (epicteatime):

hello?

OpenStudy (epicteatime):

....

OpenStudy (epicteatime):

0.559g/Cr?

OpenStudy (epicteatime):

phi could you help?

OpenStudy (jfraser):

when you divide mass by molar mass, you get units of moles. you've just found the moles of chromium. now use the balanced reaction to find the moles of chromium (II) oxide that get formed

OpenStudy (epicteatime):

thanks!!

OpenStudy (phi):

you start with 143 grams of chromium a mole of chromium weighs 52 grams \[ \frac{ 143 \text{ g}}{52 \text{ g/mole}} = 2.75 \text{ moles} \]

OpenStudy (epicteatime):

ok

OpenStudy (epicteatime):

if you are kind enough sir, could you also help (later) with another question?

OpenStudy (phi):

your equation says 4Cr + 3O2 -> 2Cr2O3 the 4 Cr must be the 2.75 moles on the other side we have 2 Cr2 O3 which is ½ the # of molecules (i.e. moles) of Cr in other words, for every 4 molecules of Cr we get 2 molecules of Cr2 O3 that means 2.75/2 moles = 1.375 moles of Cr2 O3

OpenStudy (epicteatime):

ok?

OpenStudy (phi):

4 molecules of Cr combine with 3 molecules of oxygen to give 2 molecules of chromium oxide. That means you have ½ as many molecules of chromium oxide as molecules of chromium. (4 mol of Cr become 2 mol of chromium oxide) If you started with 2.75 moles * 6.0221415 10^23 molecules / mole of chromium you will end up with half that many or 1.375 moles * 6.0221415 10^23 molecules / mole if you want that in grams, you add up how much each molecule weighs. 1 mole (that is , 6.0221415 10^23 molecules ) of \( Cr_2O_3\) weighs 52*2+16*3= 152 grams

OpenStudy (epicteatime):

THANKS DOOD